40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2456     Entry time: Mon Dec 28 10:29:31 2009
Author: Jenne 
Type: Update 
Category: Computers 
Subject: Monday Morning Bootfest 

Nothing like a good ol' Bootfest to get back into the swing of things after vacation....

It was a regular bootfest, keying crates and running everyone's startup.cmd .  There wasn't any RFM funny business which we had been dealing with a lot earlier in December (maybe Kiwamu took care of that part of things last night).

After finishing the bootfest, I tried to re-enable the watchdogs.  I noticed that the optics weren't damping at all (not that any of them were swinging crazily, they just weren't damped like regular).  This was traced to the OSEM sensor inputs and outputs being disabled on all of the suspensions' screens.  I suspect that no burt-restoring happened after c1dcuepics was powercycled yesterday. 

All of the optics are now happy as clams.

ELOG V3.1.3-